OpenStudy (anonymous):

1800m of fencing is available to fence 6 identical pig pens as shown. if the area enclosed is to be a maximum, what is the shape of each pen?

OpenStudy (anonymous):

do you have the options

OpenStudy (anonymous):

The questions are : a) explain why 9x + 8y = 1800 b)Show that the total area of each pen is given by A= -9/8x^2 + 225x m^2

OpenStudy (anonymous):

Quadratic functions

OpenStudy (anonymous):

c) if the area enclosed is to be a maximum, what is the shape of each pen?

OpenStudy (anonymous):

the shape would be the rectangle

OpenStudy (cwrw238):



OpenStudy (cwrw238):

let length of each pen be x and width be y then area of the whole = 3x * 2y length of perimeter = 3x + 3x + 2y + 2y = 6x + 4y = 1800 is this the setup? you say in for question 'as shown' but there is no diagram